What are the dimension of a 2116 square foot rectangle?
Answer by Alan3354(69443)

What are the dimension of a 2116 square foot rectangle?
-----------
There are a lot of possibilities. An infinite number, actually.
2116 = 1*2116
= 2*1058
= 4*529
= pi by 2116/pi
= 0.001 by 2116000
and many more.